OpenStudy (anonymous):

write the slope-intercept equation for the line that passes through (-7,6) and is perpendicular to -7x+9y=-2

OpenStudy (anonymous):

Using the formula of the line -7x+9y=-2 Rearrange it to, 9y= 7x-2 y= 7/9 x - 2/9 The gradient of this line is 7/9 The gradient perpendicular to this will be -9/7 Substitute (-7,6) and gradient=-9/7 to formula for a line y=mx+c 6= (-9/7)(-7) + c c = -3 Thus the equation of the line is, y=(-9/7)x -3